the number increased by 18 equal 60
Found 2 solutions by lauren_x3, bucky:
Answer by lauren_x3(3)

Step A) ?+18=60
Step B) 60-18=42
Step C) 42 increased by 18 equals 60.

Answer by bucky(2189)

The words "the number" tells you that there is some unknown number. Let's call that unknown
number "x". Next you are told that this unknown number is to be increased by 18. That means
you are to add 18 to x. In algebraic form this is written as:
.
x + 18
.
Finally, you are told that this algebraic form is to equal 60. In equation form this becomes:
.
x + 18 = 60
.
You can interpret this as some number (x) plus 18 results in 60. From algebraic procedures
we can solve for x by getting rid of the 18 on the left side of the equation. We do this
by subtracting 18 from both sides of the equation ... to get:
.
x + 18 - 18 = 60 - 18
.
Notice on the left side that the +18 and the -18 cancel each other out, so on the left
side only x remains. On the right side the 60 take away 18 results in 42. So the equation
you are left with is just:
.
x = 42
.
This is the answer to the problem. What number if you add 18 to it gives you 60? The answer
is that the number is 42 because if you add 18 to 42 you get 60.
.
Hope this helps you to understand the problem.
.